### v1

# DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators

Abstract
Generally, the decoder-only large language models (LLMs) are adapted to context-aware neural machine translation (NMT) in a concatenating way, where LLMs take the concatenation of the source sentence (i.e., intra-sentence context) and the inter-sentence context as the input, and then to generate the target tokens sequentially. This adaptation strategy, i.e., concatenation mode, considers intra-sentence and inter-sentence contexts with the same priority, despite an apparent difference between the two kinds of contexts. In this paper, we propose an alternative adaptation approach, named Decoding-enhanced Multi-phase Prompt Tuning (DeMPT), to make LLMs discriminately model and utilize the inter- and intra-sentence context and more effectively adapt LLMs to context-aware NMT. First, DeMPT divides the context-aware NMT process into three separate phases. During each phase, different continuous prompts are introduced to make LLMs discriminately model various information. Second, DeMPT employs a heuristic way to further discriminately enhance the utilization of the source-side inter- and intra-sentence information at the final decoding phase. Experiments show that our approach significantly outperforms the concatenation method, and further improves the performance of LLMs in discourse modeling.111As this paper is under review, we will release our code and datasets later.
1 Introduction
Context-aware neural machine translation (NMT) goes beyond sentence-level NMT by incorporating inter-sentence context at the document level Zhang et al. (2018); Miculicich et al. (2018); Voita et al. (2018, 2019b, 2019a); Bao et al. (2021); Sun et al. (2022), aiming to address discourse-related challenges such as zero pronoun translation Wang et al. (2019), lexical translation consistency Lyu et al. (2021, 2022), and discourse structure Hu and Wan (2023). A recent paradigm shift has been witnessed in context-aware NMT with the emergence of the decoder-only large language models (LLMs) BigScience (2022); Google (2022); MetaAI (2023b, a); OpenAI (2023). These generative language models, trained on extensive public data, have gained significant attention in the natural language processing (NLP) community. In adapting LLMs to context-aware NMT, a common strategy involves concatenating multiple source sentences as a prefix and generating translations token-by-token, relying on the prefix and previously predicted target tokens, as shown in Figure 1 (a). However, a critical observation of this strategy reveals a potential drawback – the equal prioritization of the inter- and intra-sentence contexts during token generation. Importantly, the intra-sentence context inherently contains richer parallel semantic information with the target sentence and should be given a higher priority than the inter-sentence context. Consequently, we propose that separately modeling and utilizing the inter- and intra-sentence contexts should explicitly inform LLMs of the document-level context and the current sentence itself, thus being able to prevent the misallocation of attention weights to source-side tokens Bao et al. (2021); Li et al. (2023). Inspired by the success of prompt tuning Li and Liang (2021); Liu et al. (2022); Tan et al. (2022), our alternative approach, named Decoding-Enhanced Multi-phase Prompt Tuning (DeMPT), aims to enhance LLMs’ adaptability to context-aware NMT, as shown in Figure 1 (b).
Specifically, we divide the whole procedure of context-aware NMT into three phases: inter-sentence context encoding, intra-sentence context encoding, and decoding. Following Li and Liang (2021); Liu et al. (2022), we sequentially and differentially adapt LLMs for each phase, utilizing phase-specific trainable prompts. This phased tuning method enables LLMs to independently capture and model both inter- and intra-sentence contexts, facilitating a better understanding of their differences. Importantly, our approach only divides the original input into three parts without significantly increasing computational load. As a result, there is no substantial decrease in inference speed compared to the concatenating method, as detailed in Section 4.3.
Furthermore, during the decoding phase, we propose a heuristic method to emphasize the difference between inter- and intra-sentence contexts, and avoid long-distance issue when utilizing inter-sentence context. Specifically, at each decoding step, we use LLMs to predict the next token three times. The decoding states used for each prediction directly concatenate with the representations of two contexts in a discriminative manner. Finally, we combine three probability distributions to search for the next token as the output from the target vocabulary. This method enables LLMs to learn not only to properly capture inter-sentence context in addressing discourse-related issues but also to recognize a difference between inter- and intra-sentence contexts, allowing for effective utilization of both types of contexts.
In summary, our contributions can be outlined as follows:
We propose a novel multi-phase prompt tuning approach to divide context-aware NMT into three phases, making LLMs aware of the distinction between inter- and intra-sentence contexts.
We introduce a enhanced decoding method that discriminately utilize both context types. This allows LLMs not only properly capture inter-sentence context in addressing discourse-related issues, but also be aware of the importance of the intra-sentence context.
We validate our approach using llama-2-7b and bloomz-7b1-mt as foundation models, demonstrating its effectiveness across five context-aware translation directions. Extensive analyses further highlight the substantial enhancement in LLMs’ ability for context-aware NMT.
2 Methodology
In this section, we describe our decoding-enhanced multi-phase approach for adapting LLMs to context-aware NMT in details. Specifically, we break down the whole procedure of context-aware NMT into three phases (Section 2.1), i.e., inter-sentence context encoding, intra-sentence encoding, and decoding. Additionally, we discriminatively enhance the utilization of inter- and intra-sentence contexts during the decoding phase (Section 2.2). Finally, we describe our phase-aware prompts and training objective in Section 2.3 and Section 2.4, respectively.
For a given document pair with sentences, we will construct training instances. Each training instance is denoted as a tuple . Here represents -th current source sentence with tokens, i.e., intra-sentence context, and is the -th target sentence with tokens. denotes the previous sentences of , i.e., the inter-sentence context of . We denote the hidden size of the LLM as , and as the number of transformer layers within it.
2.1 Multi-phase Encoding and Decoding
We implement our approach based on deep prompt tuning Li and Liang (2021); Liu et al. (2022). Next, we use training instance as an example to describe the multi-phase approach. Figure 2 illustrates the procedure of multi-phase prompt tuning.
Inter-sentence Context Encoding Phase.
In the inter-sentence context encoding phase (Phase 1 in Figure 2), we first concatenate all sentences in into a sequence, and then utilize the LLM to encode by incorporating the trainable prompt:
where is the sequence of activations for , is the current-phase trainable prompt, and is a hyper-parameter for the length of the prompt. aims to adapt the LLM for better modeling the inter-sentence context. Same as basic deep prompting, at the -th transformer block, we inject corresponding prompt in into encoding procedure of as follows:
where is the output of the -th transformer block. FFN and Multi-Attn are the feed-forward network sublayer and multi-head self-attention sublayer, respectively.222For simplicity, we omit the normalization and residual operations in this paper. and are the concatenating and slicing operations, respectively.
Intra-sentence Context Encoding Phase.
In the intra-sentence context encoding phase (Phase 2 in Figure 2), the LLM encodes the intra-sentence context by conditioning on the past activations of the inter-sentence context and trainable prompt:
where is the sequence of activations for , and denotes current-phase prompt. Similarly, at the -th transformer block, we incorporate and into the encoding procedure of as follows:
where is output of the -th transformer block, which fuses , the layer output of the inter-sentence context encoding.
Decoding Phase.
In the decoding phase (Phase 3 in Figure 2), given the past activations and trainable prompt, we call the LLM again to generate the hidden state for predicting the probability of the target sentence:
where is the sequence of activations for , and is current-phase prompt. Similarly, we inject and into the decoding procedure of as follows:
where is the decoding state of the -th transformer block. Finally, we refer the -th decoding state as (i.e., ) which is used to predict the next token :
where is parameter of LLM-Head layer and is the vocabulary size.
2.2 Enhanced Decoding Phase
As shown in Figure 2, both the inter-sentence context representation and the intra-sentence context representation are used as keys and values when generating hidden states of next phase. Meanwhile, hidden states of decoding phase, i.e., are used to predict next tokens. On the one hand, while the decoding hidden states incorporate both inter- and intra-sentence contexts, there is no explicit differentiation between the two when predicting next tokens. On the other hand, the inter-sentence context representation and decoding hidden states are mediated by hidden states of phases 2, i.e., . This may result in a long-distance issue such that the inter-sentence context are not properly aligned by target-side tokens.
Therefore, to address above two issues, we propose an enhanced decoding phase with an aim to more effectively utilize both the inter- and intra-sentence contexts. Inspired by Kuang et al. (2018), we move both the two types of inter- and intra-sentence contexts closer to target words to achieve a tight interaction between them. Specifically, we concatenate the decoding states with the two types of representations to predict the next target words. As shown in Figure 3, the enhanced next word prediction is a combination of three distributions with different inputs:
where and control the contribution of and , respectively, which can be further computed as:
where is same as in Eq. 16, and are the averaged and at token level, respectively.
2.3 Phase-aware Prompts
We emphasize the LLM needs to play various roles across three phases, and maintaining similar prompts across different phases may not be reasonable. Thus, we empower LLM to distinguish different phases by introducing a type embedding and a transfer layer333Different from the multi-layer perceptron (MLPs) used for reparameterization, our transfer layer is shared-parameter for all prompts. Thus, there are fewer trainable parameters during the training of our model. We compare the number of trainable parameters among different tuning methods in Table 3 and analyze the effect of the transfer layer in Appendix D. for these prompts:
where is randomly initialized prompt, are trainable parameters, and TypeEmb() is type embeddings layer of prompts. represents either phase 1, phase 2, or phase 3.
2.4 Training Objective
We employ the cross-entropy loss as the training objective of our model. Given a training instance , its training loss is defined as:
3 Experimentation
| Model | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | |
| Trans. | 29.86 | 0.8406 | 38.53 | 0.8545 | 41.44 | 0.8682 | 48.74 | 0.8783 | 32.25 | 0.8169 |
| llama-2-7b as foundation model | ||||||||||
| MT-LoRA | 27.43 | 0.8511 | 38.18 | 0.8647 | 40.96 | 0.8712 | 47.52 | 0.8733 | 33.00 | 0.8311 |
| MT-PT | 31.32 | 0.8565 | 41.92 | 0.8675 | 43.56 | 0.8752 | 51.32 | 0.8819 | 35.46 | 0.8333 |
| CMT-PT | 31.13 | 0.8387 | 42.01 | 0.8699 | 43.11 | 0.8762 | 51.66 | 0.8823 | 35.91 | 0.8396 |
| MPT | *33.21 | 0.8645 | †43.11 | 0.8744 | *43.88 | 0.8824 | †52.01 | 0.8913 | †36.49 | 0.8456 |
| DeMPT | *33.89 | 0.8658 | †43.71 | 0.8816 | *44.69 | 0.8899 | †53.10 | 0.8979 | †36.55 | 0.8438 |
| bloomz-7b1-mt as foundation model | ||||||||||
| MT-LoRA | 25.79 | 0.8466 | 35.67 | 0.8601 | 35.17 | 0.8522 | 46.32 | 0.8644 | 28.01 | 0.8012 |
| MT-PT | 30.99 | 0.8520 | 40.49 | 0.8661 | 37.76 | 0.8579 | 50.68 | 0.8823 | 30.27 | 0.8106 |
| CMT-PT | 30.82 | 0.8504 | 40.31 | 0.8639 | 38.01 | 0.8601 | 50.26 | 0.8832 | 29.80 | 0.8108 |
| MPT | *31.81 | 0.8601 | *41.11 | 0.8766 | †38.99 | 0.8669 | *51.33 | 0.8910 | *30.99 | 0.8201 |
| DeMPT | *32.46 | 0.8649 | *41.92 | 0.8790 | †40.06 | 0.8703 | *52.25 | 0.8990 | *31.79 | 0.8253 |
| Model | ZH | FR | DE | ES | RU |
| Trans. | 47.63 | 54.41 | 58.29 | 62.52 | 48.79 |
| llama-2-7b as foundation model | |||||
| MT-LoRA | 44.83 | 54.52 | 57.72 | 62.18 | 49.06 |
| MT-PT | 49.49 | 57.87 | 60.89 | 65.02 | 52.59 |
| CMT-PT | 49.53 | 58.27 | 61.23 | 65.89 | 53.34 |
| MPT | 51.56 | 59.56 | 62.15 | 67.14 | 54.18 |
| DeMPT | 52.68 | 60.33 | 63.11 | 67.95 | 54.34 |
| bloomz-7b1-mt as foundation model | |||||
| MT-LoRA | 43.23 | 51.82 | 51.12 | 61.77 | 43.29 |
| MT-PT | 49.48 | 56.81 | 55.40 | 64.71 | 46.14 |
| CMT-PT | 49.61 | 57.05 | 55.81 | 65.12 | 46.09 |
| MPT | 50.22 | 57.93 | 56.69 | 66.25 | 47.29 |
| DeMPT | 50.62 | 58.30 | 57.34 | 67.12 | 48.00 |
We build our approach upon two open-source LLMs, namely, llama-2-7b444https://huggingface.co/meta-llama/Llama-2-7b-hf and bloomz-7b1-mt555https://huggingface.co/bigscience/bloomz-7b1-mt. We verify the effectiveness of our proposed approach on five translation tasks, including {Chinese (ZH), French (FR), German (DE), Spanish (ES), Russian (RU)}English (EN).
3.1 Experimental Settings
Datasets and Preprocessing.
The corpus of all translation tasks is extracted from New-Comentary-v18. See Appendix A for splitting and statistics of the training set, valid set, and test set. We use the tokenizer of foundation models to process the input data and no any other preprocessing is performed.
Baselines.
We compare our approach against four baselines:
Transformer (Trans.): It is an encoder-decoder Transformer-base model Vaswani et al. (2017) that is trained from scratch.
MT-LoRA: It is a tuned LLM adapted to NMT task via the tuning method of Low-Rank Adaptation Hu et al. (2022), which makes large-scale pre-training models adapt to a new task by injecting a trainable rank decomposition matrice into each layer of the Transformer architecture.
MT-PT: It is a tuned LLM adapted to NMT task via the deep prompt tuning with MLPs reparameterization,666We attempt to remove reparameterization but experience a significant decline in performance. which only tunes continuous prompts with a frozen language model.
CMT-PT: Similar to MT-PT, it is also a tuned LLM via the deep prompt tuning with MLPs reparameterization. Unlike MT-PT, it utilizes inter-sentence context within the concatenation strategy, as depicted in Figure 1 (a).
Among them, Transformer, MT-LoRA, and MT-PT are context-agnostic systems while CMT-PT is a context-aware system. For a fair comparison, we ensure that all context-aware systems, including CMT-PT, MPT, and DeMPT, incorporate identical inter-sentence context.
Model Setting and Training.
For the Transformer model, we implement it upon Fairseq Ott et al. (2019). For MT-LoRA models, we set the rank of trainable matrices as 16 which performs best in our preliminary experiment. For all MT-PT models, CMT-PT models, and our models, we set the prompt length as 64. For the incorporation of inter-sentence context in CMT-PT models and our models, we consider a dynamic , in which the total tokens are no more than 256. In enhanced decoding, we consider the three next word predictions to be equally important by setting both and to 1/3. More details of training are provided in Appendix B.
Evaluation.
We use sacreBLEU (accuracy-related metric)777Signature: nrefs:1|case:mixed|eff:no|tok:13a|
smooth:exp|version:2.3.1 Post (2018), COMET (semantics-related metric) with the wmt22-comet-da model888https://github.com/Unbabel/COMET Rei et al. (2020), and BlonDe (discourse-related metric) Jiang et al. (2022) as the evaluation metrics.
3.2 Experimental Results
The main experimental results are presented in Tables 1 and 2. Additionally, a comparison of the number of trainable parameters is presented in Table 3 across different tuning methods. From these results we have the following observations:
The encoder-decoder Transformer model (Trans.) performs better than the LLMs with LoRA tuning in most translation directions in BLEU score. For example, when utilizing llama-2-7b as the foundation model, Transformer surpasses MT-LoRA an average of 0.75 BLEU score across all translation tasks. However, MT-LoRA model outperforms Trans. in terms of COMET, suggesting that translations from LLMs may align more closely with human preferences.
The MT-PT model presents superior performance compared to the MT-LoRA model in terms of BLEU, COMET, and BlonDe. Taking bloomz-7b1-mt as the foundation model, the MT-PT model outperforms the MT-LoRA model by an average of 3.84 BLEU score, 3.51 BlonDe score, and 0.0047 COMET score. Nevertheless, the MT-PT model sacrifices efficiency for performance, introducing more trainable parameters (13.87% vs. 0.12%).
Leveraging the inter-sentence context is helpful in alleviating discourse-related issues. For example, with bloomz-7b1-mt used as the foundation model, the CMT-PT model, despite underperforming in BLEU and COMET compared to the MT-PT model, excels in discourse-related BlonDe scores (averaging 57.66 vs. 57.17).
Our MPT/DeMPT model outperforms all baselines across all translation tasks. For example, when using llama-2-7b as the foundation model, our MPT model achieves an average gain of 1.62/1.45/2.03 in BLEU/COMET/BlonDe score compared to the CMT-PT model. Furthermore, our decoding-enhance strategy enhances the capacity of LLMs in context-aware NMT, with DeMPT outperforming MPT in BLEU/COMET/BlonDe score (averaging 42.39/0.8758/59.68 vs. 41.74/0.8716/58.91).
The model built upon llama-2-7b as the foundation model outperforms the one using bloomz-7b1-mt, suggesting that llama-2-7b serves as a more robust foundation model for translation tasks.
| MT-LoRA | MT-PT/CMT-PT | DeMPT | |
| Trainable Para. | 0.12% | 13.87% | 3.11% |
4 Discussion
In this section, we use bloomz-7b1-mt as the foundation model to discuss and analyze our approach. See Appendix CE for further discussions.
4.1 Effect of Length of Inter-sentence Context
For efficient training, we define the inter-sentence context in Section 2 as previous sentences with a total tokens not exceeding 256. We are curious about the potential impact of inter-sentence length on the performance of our approach. Consequently, we extend the inter-sentence context length from 256 to 1024 and assess the performance of our approach in the ZHEN task.
Figure 4 shows the performance trend of the CMT-PT model and our DeMPT model. As the length of the inter-sentence context increases, both models exhibit a slight enhancement in both BLEU and BlonDe scores. Interestingly, our model with a 256-token inter-sentence context outperforms the CMT-PT model with a 1024-token inter-sentence context in both BLEU and BlonDe scores. This further suggests the effectiveness of our approach in harnessing the capabilities of LLMs for context-aware NMT compared to the concatenation strategy.
4.2 Effect of Prompt Length
As our approach is implemented based on deep prompt tuning, next we compare the impact of the trainable prompt length for MT-PT, CMT-PT, and our DeMPT.
Figure 5 shows the performance curves when increasing the prompt length from 32 to 128. We observe that increased prompt length tends to enhance performance for both BLEU and BlonDe, yet the gains exhibit diminishing returns. This finding is consistent with that in Li and Liang (2021); Lester et al. (2021); Tan et al. (2022). We also observe that DeMPT with a prompt length of 64 outperforms both MT-PT and CMT-PT with a prompt length of 128 on both metrics, suggesting the superiority of our approach over the concatenation strategy in enhancing LLMs’ capacity for context-aware NMT.
4.3 Comparison of Inference Speed
| Model | Speed | BLEU |
| MT-PT | 0.75 sec/sent. | 30.99 |
| CMT-PT | 0.77 sec/sent. | 30.82 |
| MPT | 0.78 sec/sent. | 31.81 |
| DeMPT | 0.79 sec/sent. | 32.46 |
Table 4 compares the inference speed of different models on ZHEN translation task. Our MPT and DeMPT models, dividing the context-aware NMT process into three separate phases, demonstrates comparable inference speed to the single-phase MT-PT and CMT-PT models, with only a marginal drop of 0.02 seconds per sentence in decoding. This illustrates the efficiency of our approach without introducing significant computational overhead.
4.4 Performance on Contrastive Test Set
We evaluate the models’ ability to resolve discourse inconsistencies using the contrastive test set proposed by Voita et al. (2019a), which focuses on four discourse phenomena such as deixis, lexicon consistency (lex.c), ellipsis inflection (ell.infl), and verb phrase ellipsis (ell.VP) in EnglishRussian translation. Within the test set, each instance comprises a positive translation and several negative ones that vary by only one specific word. The purpose of the contrastive test set is to assess whether a model is more inclined to generate a correct translation as opposed to incorrect variations.
Table 5 lists the accuracy of translation prediction on the contrastive test set for MT-PT, CMT-PT and DeMPT. Compared to the context-agnostic MT-PT model, both context-aware CMT-PT and DeMPT models show substantial improvements across the four discourse phenomena. Additionally, DeMPT demonstrates the best performance, surpassing CMT-PT by an average accuracy margin of 3.8.
| Model | deixis | lex.c | ell.infl | ell.VP | Avg. |
| MT-PT | 50.0 | 45.7 | 53.0 | 28.6 | 44.3 |
| CMT-PT | 80.2 | 46.1 | 74.3 | 75.3 | 68.9 |
| DeMPT | 80.1 | 55.7 | 75.9 | 79.3 | 72.7 |
4.5 Human Evaluation
We use the Direct Assessment (DA) method Graham et al. (2017) to manually assess the quality of translations generated by DeMPT and CMT-PT. In this assessment, human evaluators compare the meaning of the MT output with a human-produced reference translation, working within the same language.
Specifically, we randomly select 5 documents with a total of 200 groups of sentences from the ZHEN test set. To avoid potential bias in evaluation, we recruit 6 professional translators and ensure each translation from DeMPT or CMT-PT is scored twice by two translators. Table 6 shows the DA scores for CMT-PT and DeMPT. Our DeMPT outperforms CMT-PT by 7.14 DA score, providing strong evidence for the effectiveness of our approach. Further details and results regarding the DA can be found in Appendix C.
| Model | Score_1 | Score_2 | Average |
| CMT-PT | 79.00 | 80.17 | 79.59 |
| DeMPT | 86.17 (+7.17) | 87.30 (+7.13) | 86.73 (+7.14) |
5 Related Work
Large Language Models for Context-aware Machine Translation.
While traditional context-aware neural machine translation (NMT) has seen considerable progress in recent years Jean et al. (2017); Wang et al. (2017); Voita et al. (2018); Maruf et al. (2019); Kang et al. (2020); Bao et al. (2021); Sun et al. (2022); Bao et al. (2023), the effective integration of large language models (LLMs) to model inter-sentence context and enhance context-aware translation remains an area of limited exploration. Existing studies mainly focus on the assessment of LLMs’ ability in discourse modeling. For example, Wang et al. (2023) approach context-aware NMT as a task involving long sequence generation, employing a concatenation strategy, and conduct comprehensive evaluations of LLMs such as ChatGPT and GPT-4. Their focus includes the impact of context-aware prompts, comparisons with translation models, and an in-depth analysis of discourse modeling ability. Similarly, Karpinska and Iyyer (2023) engage professional translators to evaluate LLMs’ capacity in context-aware NMT. In contrast, Wu et al. (2024) compare the effectiveness of various parameter-efficient fine-tuning methods on moderately-sized LLMs for context-aware NMT. Besides, Wu and Hu (2023) explore the prompt engineering with GPT language models specifically for document-level (context-aware) MT while Li et al. (2024) experiment with combining sentence-level and document-level translation instructions of varying lengths to fine-tune LLMs.
Prompt Tuning for Large Language Model.
Liu et al. (2021) and Li and Liang (2021) propose to make LLMs adapt to various tasks by adding trainable prompts (also called continuous prompts) to the original input sequences. In this paradigm, only the continuous prompts are updated during training. Liu et al. (2022) further introduce deep prompt tuning, extending the idea by inserting trainable prompts into all layers of LLMs, rather than just the embedding layer. While these approaches lay the groundwork for a general framework, our focus lies in augmenting the performance of LLMs specifically for inter-sentence context modeling in context-aware NMT. Notably related, Tan et al. (2022) propose a multi-phase tuning approach to enhance the sentence-level translation performance of a multilingual GPT. However, the exploration of effective LLM tuning for addressing discourse-related challenges in the context-aware NMT domain remains underdeveloped.
6 Conclusion
In this paper, we have examined the hypothesis that it is crucial to differentially model and leverage inter-sentence context and intra-sentence context when adapting LLMs to context-aware NMT. This stems from our observation that intra-sentence context exhibits a stronger correlation with the target sentence compared to inter-sentence context, owing to its richer parallel semantic information. To this end, we have proposed a novel decoding-enhanced multi-phase prompt tuning (DeMPT) approach to make LLMs aware of the differences between inter- and intra-sentence contexts, and further improve LLMs’ capacity in discourse modeling. We have evaluated our approach using two foundation models and present experimental results across five translation directions. Experimental results and discussions have demonstrated a significant enhancement in the performance of LLMs in context-aware NMT, manifesting as improved translation accuracy and a reduction in discourse-related issues.
Limitations
Owing to resource limitations, our work is restricted to moderate-scale LLMs, specifically those with 7 billion parameters, and a confined window size of inter-sentence context. It is imperative to acknowledge that the results of our research may differ when employing larger models and an extended window size for inter-sentence context. We acknowledge these limitations and consider them as avenues for future exploration.
Appendix A Datasets
Statistics and Splitting of Datasets.
We provide the detailed statistic in Table 7. For all translation tasks, we randomly select 80% document pairs from the corpus as the training set. Both the test set and validation set include 150 document pairs each, randomly sampled from the remaining 20% of document pairs in the corpus. Regarding sentence preprocessing across all datasets, we segment the sentences with the tokenizer from the respective foundation model. No additional preprocessing steps are performed. Datasets are downloaded from https://data.statmt.org/news-commentary/v18.
| Dataset | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | |
| Training | 8,622 | 342,495 | 7,915 | 310,489 | 8,417 | 333,201 | 9,677 | 378,281 | 7,255 | 272,100 |
| Validation | 150 | 6,061 | 150 | 5,890 | 150 | 5,866 | 150 | 5,782 | 150 | 5,691 |
| Test | 150 | 5,747 | 150 | 5,795 | 150 | 5,967 | 150 | 5,819 | 150 | 5,619 |
Appendix B Training Details
For all Transformer NMT models, we use the transformer-base setting as in Vaswani et al. (2017), where the learning rate is set to 1e-4. The Transformer NMT models are trained on 4 NVIDIA V100 32GB GPUs with a batch size of 4096. For the models with prompt tuning in Section 3, including MT-PT, CMT-PT, and our MPT and DeMPT models, the length of the trainable prompt is set as 64. During both training and inference, the model generates only the current target sentence, operating in a many-to-one translation mode. For all fine-tuning models in this paper, we set the training epoch to 4, and the warm-up rate to 0.1. We use the log learning rate decay strategy with a maximum learning rate of 5e-5. We collate a mini-batch by counting the total tokens inside the batch and set the batch size as 4096. All fine-tuning models are trained on 4 NVIDIA A800 GPUs with Deespeed Zero 2 offload setting Rajbhandari et al. (2020).999https://github.com/microsoft/DeepSpeed
Appendix C Details of Human Evaluation
Criterion and Recruitment.
Given a source sentence, its translation from MT (i.e., CMT-PT and our DeMPT), and its human-produced reference translation, the evaluators are asked to give a score ranging from 0 to 100. Figure 6 presents the detailed criterion of scoring. We recruit evaluators from professional translators with at least five years of experience in translation.
Statistics of Translation Errors.
We manually count the number of bad cases from our DeMPT model. The bad cases fall into two categories: (1) the DA score is 60 or lower; (2) the DA score is lower than that of the translation from CMT-PT. The main types of the bad cases are Mistranslation (Mis.), Unnoticed Omission (UO), Inappropriate Expression (IE), and Grammatical Error (GE). We present detailed statistics in Table 8. The statistics indicate the bad cases mainly come from Mistranslation and Unnoticed Omission. Meanwhile, our DeMPT model outperforms the CMT-PT model in 86.5% DA cases.
Case Study.
We present a case in Figure 7 to illustrate how our DeMPT model outperforms the CMT-PT model. In this case, we compare the translations of two consecutive sentences from our model and the CMT-PT model. First, we notice that the CMT-PT model translates the source word 美国 in the two sentences into US and America, respectively. However, our model consistently translates them into US. Second, our model uses for its part, a phase with more coherent preference, as the translation of 同时 , instead of At the same time adopted in the translation from the CMT-PT model. Both of them demonstrate the superiority of our proposed approach in discourse modeling.
| Group | Type of Bad Case | ||||
| Mis. | UO | IE | GE | Total (Perc.) | |
| 1 | 6 | 3 | 1 | 2 | 12 (6.0%) |
| 2 | 9 | 7 | 6 | 5 | 27 (13.5%) |
Appendix D Effect of Transfer Layer and Type Embedding
As in Eq. 22 within Section 2.3, we introduce two sublayers: a non-linear transfer sublayer and a type embedding sublayer for the trainable prompt in each phase. This design enhances the awareness of LLMs regarding the distinctions in inputs across the three tuning phases, allowing them to adapt to specific roles at each phase. We investigate the effect of these two sublayers.
As shown in Table 9, our observations reveal that the transfer sublayer holds greater importance than the type embedding sublayer. Removing either the non-linear transfer sublayer (w/o Transfer.) or the type embedding sublayer (w/o Embed.) results in a performance drop of 0.84/0.0048/0.39 or 0.45/0.0036/0.007 in BLEU/COMET/BlonDe metrics.
| Model | BLEU | COMET | BlonDe |
| MT-PT | 30.99 | 0.8520 | 49.48 |
| CMT-PT | 30.82 | 0.8504 | 49.61 |
[…]
| w/o Transfer. | 31.62 | 0.8601 | 50.23 |
| w/o Embed. | 32.01 | 0.8613 | 50.55 |
| w/o CTX. | 31.98 | 0.8593 | 49.89 |
Appendix E Effect of Inter-sentence Context
We implement the context-agnostic (sentence-level) DeMPT system to analyze the effect of the inter-sentence context. More specifically, we replace the input of LLMs in the inter-sentence context encoding phase with the intra-sentence context. In other words, we encode the intra-sentence context twice to keep the multi-phase tuning strategy in DeMPT while making the inter-sentence context unavailable.
As shown in the last row of Table 9 (i.e., w/o CTX), we find that the inter-sentence context is crucial for the alleviation of discourse-related issues. The BlonDe score drops by 0.73 when the inter-sentence context is unavailable. Meanwhile, our DeMPT also significantly improves the performance of LLMs in context-agnostic MT, e.g., + 0.99 BLEU score and + 0.0073 COMET score compared to the MT-PT model.