### v2

# DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators

Abstract
Generally, the decoder-only large language models (LLMs) are adapted to context-aware neural machine translation (NMT) in a concatenating way, where LLMs take the concatenation of the source sentence (i.e., intra-sentence context) and the inter-sentence context as the input, and then to generate the target tokens sequentially. This adaptation strategy, i.e., concatenation mode, considers intra-sentence and inter-sentence contexts with the same priority, despite an apparent difference between the two kinds of contexts. In this paper, we propose an alternative adaptation approach, named Decoding-enhanced Multi-phase Prompt Tuning (DeMPT), to make LLMs discriminately model and utilize the inter- and intra-sentence context and more effectively adapt LLMs to context-aware NMT. First, DeMPT divides the context-aware NMT process into three separate phases. During each phase, different continuous prompts are introduced to make LLMs discriminately model various information. Second, DeMPT employs a heuristic way to further discriminately enhance the utilization of the source-side inter- and intra-sentence information at the final decoding phase. Experiments show that our approach significantly outperforms the concatenation method, and further improves the performance of LLMs in discourse modeling.111We release our code and datasets on https://github.com/xllyu-nlp/DeMPT.
DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators
Xinglin Lyu♣, Junhui Li♣††thanks: Corresponding author: Junhui Li., Yanqing Zhao♠, Min Zhang♠, Daimeng Wei♠, Shimin Tao♠, Hao Yang♠, Min Zhang♣ ♣School of Computer Science and Technology, Soochow University, Suzhou, China ♠Huawei Translation Services Center, Beijing, China xllv2020@stu.suda.edu.cn, {lijunhui,minzhang}@suda.edu.cn {zhaoyanqing,zhangmin186,weidaimeng,taoshimin,yanghao30}@huawei.com
1 Introduction
Context-aware neural machine translation (NMT) goes beyond sentence-level NMT by incorporating inter-sentence context at the document level Zhang et al. (2018); Miculicich et al. (2018); Voita et al. (2018, 2019b, 2019a); Bao et al. (2021); Sun et al. (2022), aiming to address discourse-related challenges such as zero pronoun translation Wang et al. (2019), lexical translation consistency Lyu et al. (2021, 2022), and discourse structure Hu and Wan (2023). A recent paradigm shift has been witnessed in context-aware NMT with the emergence of the decoder-only large language models (LLMs) BigScience (2022); Google (2022); MetaAI (2023b, a); OpenAI (2023). These generative language models, trained on massive data, have gained significant attention in the natural language processing (NLP) community. In adapting LLMs to context-aware NMT, a common strategy involves concatenating multiple source sentences as a prefix and generating translations token-by-token, relying on the prefix and previously predicted target tokens, as shown in Figure 1 (a). However, a critical observation of this strategy reveals a potential drawback – the equal prioritization of the inter- and intra-sentence contexts during token generation. Importantly, the intra-sentence context inherently contains richer parallel semantic information with the target sentence and should be given a higher priority than the inter-sentence context. Consequently, we propose that separately modeling and utilizing the inter- and intra-sentence contexts should explicitly inform LLMs of the document-level context and the current sentence itself, thus being able to prevent the misallocation of attention weights to source-side tokens Bao et al. (2021); Li et al. (2023). Inspired by the success of prompt tuning Li and Liang (2021); Liu et al. (2022); Tan et al. (2022), our alternative approach, named Decoding-Enhanced Multi-phase Prompt Tuning (DeMPT), aims to enhance LLMs’ adaptability to context-aware NMT, as shown in Figure 1 (b).222Following the findings of Bao et al. (2021), which indicate that source-side context is relatively more important for document-level MT compared to target-side context, we focus exclusively on source-side context in this paper. Nonetheless, we provide an additional discussion on integrating target-side context in Appendix M.
Specifically, we divide the whole procedure of context-aware NMT into three phases: inter-sentence context encoding, intra-sentence context encoding, and decoding. Following Li and Liang (2021); Liu et al. (2022), we sequentially and differentially adapt LLMs for each phase, utilizing phase-specific trainable prompts. This phased tuning method enables LLMs to independently capture and model both inter- and intra-sentence contexts, facilitating a better understanding of their differences. Our approach splits the input into three parts without significantly increasing computational load, thus maintaining inference speed comparable to concatenation, as detailed in Appendix D.
Furthermore, during the decoding phase, we propose a heuristic method to emphasize the difference between inter- and intra-sentence contexts, and avoid long-distance issue when utilizing inter-sentence context. Specifically, at each decoding step, we use LLMs to predict the next token three times. The decoding states used for each prediction directly concatenate with the representations of two contexts in a discriminative manner. Finally, we combine three probability distributions to search for the next token as the output from the target vocabulary. This method enables LLMs to learn not only to properly capture inter-sentence context in addressing discourse-related issues but also to recognize a difference between inter- and intra-sentence contexts, allowing for effective utilization of both types of contexts.
Our contributions can be summarized as follows:
We introduce a multi-phase prompt tuning approach that divides context-aware NMT into three phases, enabling LLMs to distinguish between inter- and intra-sentence contexts.
We introduce a enhanced decoding method that discriminately utilize both context types. This allows LLMs not only properly capture inter-sentence context in addressing discourse-related issues, but also be aware of the importance of the intra-sentence context.
We validate our approach using llama-2-7b and bloomz-7b1-mt as foundation models, demonstrating its effectiveness across five translation directions. Extensive analyses further highlight the substantial enhancement in LLMs’ ability for context-aware MT.
2 Methodology
In this section, we describe our decoding-enhanced multi-phase approach for adapting LLMs to context-aware NMT in details. Specifically, we break down the whole procedure of context-aware NMT into three phases (Section 2.1), i.e., inter-sentence context encoding, intra-sentence encoding, and decoding. Additionally, we discriminatively enhance the utilization of inter- and intra-sentence contexts during the decoding phase (Section 2.2). Finally, we describe our phase-aware prompts and training objective in Section 2.3 and Section 2.4, respectively.
For a given document pair with sentences, we will construct training instances. Each training instance is denoted as a tuple . Here represents -th current source sentence with tokens, i.e., intra-sentence context, and is the -th target sentence with tokens. denotes the previous sentences of , i.e., the inter-sentence context of . We denote the hidden size of the LLM as , and as the number of transformer layers within it.
2.1 Multi-phase Encoding and Decoding
We implement our approach based on deep prompt tuning Li and Liang (2021); Liu et al. (2022). Next, we use training instance as an example to describe the multi-phase approach. Figure 2 illustrates the procedure of multi-phase prompt tuning.
Inter-sentence Context Encoding Phase.
In the inter-sentence context encoding phase (Phase 1 in Figure 2), we first concatenate all sentences in into a sequence, and then utilize the LLM to encode by incorporating the trainable prompt:
where is the sequence of activations for , is the current-phase trainable prompt, and is a hyper-parameter for the length of the prompt. aims to adapt the LLM for better modeling the inter-sentence context. Same as basic deep prompting, at the -th transformer block, we inject corresponding prompt in into encoding procedure of as follows:
where is the output of the -th transformer block. FFN and Multi-Attn are the feed-forward network and multi-head self-attention sublayers, respectively.333For simplicity, we omit the normalization and residual operations in this paper. and are the concatenating and slicing operations, respectively.
Intra-sentence Context Encoding Phase.
In the intra-sentence context encoding phase (Phase 2 in Figure 2), the LLM encodes the intra-sentence context by conditioning on the past activations of the inter-sentence context and trainable prompt:
where is the sequence of activations for , and denotes current-phase prompt. Similarly, at the -th transformer block, we incorporate and into the encoding procedure of as follows:
where is output of the -th transformer block, which fuses , the layer output of the inter-sentence context encoding.
Decoding Phase.
In the decoding phase (Phase 3 in Figure 2), given the past activations and trainable prompt, we call the LLM again to generate the hidden state for predicting the probability of the target sentence:
where is the sequence of activations for , and is current-phase prompt. Similarly, we inject and into the decoding procedure of as follows:
where is the decoding state of the -th transformer block. Finally, we refer the -th decoding state as (i.e., ) which is used to predict the next token :
where is parameter of LLM-Head layer and is the vocabulary size.
2.2 Enhanced Decoding Phase
As shown in Figure 2, both the inter-sentence context representation and the intra-sentence context representation are used as keys and values when generating hidden states of next phase. Meanwhile, hidden states of decoding phase, i.e., are used to predict next tokens. On the one hand, while the decoding hidden states incorporate both inter- and intra-sentence contexts, there is no explicit differentiation between the two when predicting next tokens. On the other hand, the inter-sentence context representation and decoding hidden states are mediated by hidden states of phases 2, i.e., . This may result in a long-distance issue such that the inter-sentence context are not properly aligned by target-side tokens.
Therefore, to address above two issues, we propose an enhanced decoding phase with an aim to more effectively utilize both the inter- and intra-sentence contexts. Inspired by Kuang et al. (2018), we move both the two types of inter- and intra-sentence contexts closer to target words to achieve a tight interaction between them. Specifically, we concatenate the decoding states with the two types of representations to predict the next target words. As shown in Figure 3, the enhanced next word prediction is a combination of three distributions:
where and control the contribution of and , respectively, which can be further formulated as:
where is same as in Eq. 16, and are the averaged and at token level, respectively.444Notably, the computation of and does not require a full decoding forward pass. It involves solely an FFN layer (two linear transformation layers and a ReLU activation layer), an LLM-Head layer (a linear transformation layer), and a softmax function layer. To further identify the effect of inter- and intra-sentence context in this strategy, we provide an ablation study about and in Appendix J.
2.3 Phase-aware Prompts
We emphasize the LLM needs to play various roles across three phases, and maintaining similar prompts across different phases may not be reasonable. Thus, we empower LLM to distinguish different phases by introducing a type embedding and a transfer layer555Unlike the multi-layer perceptrons (MLPs) used for reparameterization, our transfer layer shares parameters across all prompts, reducing the number of trainable parameters. Table 3 compares the trainable parameters of various tuning methods, and Appendix J analyzes the effect of the transfer layer. for these prompts:
where is randomly initialized prompt, are the trainable parameters, and TypeEmb() is type embeddings layer of the prompts. represents either phase 1, phase 2, or phase 3.
2.4 Training Objective
We employ the cross-entropy loss as the training objective of our model. Given a training instance , its training loss is defined as:
3 Experimentation
| Model | ZHEN | FREN | DEEN | ESEN | RUEN | Average | ||||||
| BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | |
| ⊘Trans. | 29.86 | 0.8406 | 38.53 | 0.8545 | 41.44 | 0.8682 | 48.74 | 0.8783 | 32.25 | 0.8169 | 38.16 | 0.8517 |
| Traditional context-aware NMT models | ||||||||||||
| ⊙MR-Trans. | 30.61 | 0.8413 | 38.72 | 0.8533 | 42.11 | 0.8693 | 49.69 | 0.8812 | 33.27 | 0.8211 | 38.88 | 0.8532 |
| + mBART | 32.69 | 0.8601 | 42.01 | 0.8759 | 44.61 | 0.8840 | 51.67 | 0.8831 | 36.39 | 0.8459 | 41.39 | 0.8698 |
| ⊙G-Trans. | 30.99 | 0.8411 | 38.96 | 0.8524 | 42.46 | 0.8658 | 49.68 | 0.8794 | 33.59 | 0.8201 | 39.14 | 0.8518 |
| + mBART | 32.99 | 0.8597 | 42.02 | 0.8764 | 44.81 | 0.8836 | 52.07 | 0.8911 | 36.83 | 0.8461 | 41.74 | 0.8714 |
| llama-2-7b as foundation model | ||||||||||||
| ⊘MT-LoRA | 27.43 | 0.8511 | 38.18 | 0.8647 | 40.96 | 0.8712 | 47.52 | 0.8733 | 33.00 | 0.8311 | 37.42 | 0.8583 |
| ⊘MT-PT | 31.32 | 0.8565 | 41.92 | 0.8675 | 43.56 | 0.8752 | 51.32 | 0.8819 | 35.46 | 0.8333 | 40.72 | 0.8629 |
| ⊙CMT-PT | 31.13 | 0.8387 | 42.01 | 0.8699 | 43.11 | 0.8762 | 51.66 | 0.8823 | 35.91 | 0.8396 | 40.76 | 0.8613 |
| ⊙MPT | *33.21 | 0.8645 | †43.11 | 0.8744 | *43.88 | 0.8824 | †52.01 | 0.8913 | †36.49 | 0.8456 | 41.74 | 0.8716 |
| ⊙DeMPT | *33.89 | 0.8658 | †43.71 | 0.8816 | *44.69 | 0.8899 | †53.10 | 0.8979 | †36.55 | 0.8438 | 42.39 | 0.8758 |
| bloomz-7b1-mt as foundation model | ||||||||||||
| ⊘MT-LoRA | 25.79 | 0.8466 | 35.67 | 0.8601 | 35.17 | 0.8522 | 46.32 | 0.8644 | 28.01 | 0.8012 | 34.21 | 0.8449 |
| ⊘MT-PT | 30.99 | 0.8520 | 40.49 | 0.8661 | 37.76 | 0.8579 | 50.68 | 0.8823 | 30.27 | 0.8106 | 38.04 | 0.8539 |
| ⊙CMT-PT | 30.82 | 0.8504 | 40.31 | 0.8639 | 38.01 | 0.8601 | 50.26 | 0.8832 | 29.80 | 0.8108 | 37.84 | 0.8537 |
| ⊙MPT | *31.81 | 0.8601 | *41.11 | 0.8766 | †38.99 | 0.8669 | *51.33 | 0.8910 | *30.99 | 0.8201 | 38.85 | 0.8629 |
| ⊙DeMPT | *32.46 | 0.8649 | *41.92 | 0.8790 | †40.06 | 0.8703 | *52.25 | 0.8990 | *31.79 | 0.8253 | 39.70 | 0.8677 |
| Model | ZH | FR | DE | ES | RU | Avg. |
| ⊘Trans. | 47.63 | 54.41 | 58.29 | 62.52 | 48.79 | 54.33 |
| Traditional context-aware NMT model | ||||||
| ⊙MR-Trans. | 48.51 | 55.55 | 59.02 | 63.51 | 49.88 | 55.29 |
| + mBART | 50.66 | 58.01 | 61.99 | 66.01 | 54.11 | 58.15 |
| ⊙G-Trans. | 48.99 | 55.31 | 59.23 | 63.99 | 50.09 | 55.52 |
| + mBART | 50.98 | 57.88 | 61.97 | 66.21 | 54.33 | 58.27 |
| llama-2-7b as foundation model | ||||||
| ⊘MT-LoRA | 44.83 | 54.52 | 57.72 | 62.18 | 49.06 | 53.66 |
| ⊘MT-PT | 49.49 | 57.87 | 60.89 | 65.02 | 52.59 | 57.17 |
| ⊙CMT-PT | 49.53 | 58.27 | 61.23 | 65.89 | 53.34 | 57.65 |
| ⊙MPT | 51.56 | 59.56 | 62.15 | 67.14 | 54.18 | 58.92 |
| ⊙DeMPT | 52.68 | 60.33 | 63.11 | 67.95 | 54.94 | 59.80 |
| bloomz-7b1-mt as foundation model | ||||||
| ⊘MT-LoRA | 43.23 | 51.82 | 51.12 | 61.77 | 43.29 | 50.25 |
| ⊘MT-PT | 49.48 | 56.81 | 55.40 | 64.71 | 46.14 | 54.51 |
| ⊙CMT-PT | 49.61 | 57.05 | 55.81 | 65.12 | 46.09 | 54.74 |
| ⊙MPT | 50.22 | 57.93 | 56.69 | 66.25 | 47.29 | 55.68 |
| ⊙DeMPT | 50.62 | 58.30 | 57.34 | 67.12 | 48.00 | 56.28 |
We build our approach upon two open-source LLMs, i.e., llama-2-7b666https://huggingface.co/meta-llama/Llama-2-7b-hf and bloomz-7b1-mt777https://huggingface.co/bigscience/bloomz-7b1-mt. We verify the effectiveness of our proposed approach on five translation tasks, including {Chinese (ZH), French (FR), German (DE), Spanish (ES), Russian (RU)}English (EN).
3.1 Experimental Settings
Datasets and Preprocessing.
The corpus of all translation tasks is extracted from News-Commentary-v18. For LLM-based models, We use the tokenizer of foundation models to process the input data and no other preprocessing is performed. See Appendix A for more details on splitting, preprocessing and statistics of datasets. Besides, we provide a discussion for scales of the training set in Appendix I .
Baselines.
In addition to traditional context-agnostic (Trans.) and context-aware (G-Trans Bao et al. (2021) and MR-Trans Sun et al. (2022) with or without pre-training setting, i.e., + mBART liu2020multilingual) NMT models with encoder-decoder architecture ,888Please refer to Appendix C for more introduction about the G-Trans and MR-Trans. our primary comparison focuses on the following three LLM-based alternatives: 1) MT-LoRA: It is a tuned LLM adapted to NMT task via the tuning method of Low-Rank Adaptation Hu et al. (2022), which makes large-scale pre-training models adapt to a new task by injecting a trainable rank decomposition matrice into each layer of the Transformer architecture; 2) MT-PT: It is a tuned LLM adapted to NMT task via the deep prompt tuning with MLPs reparameterization,999We attempt to remove reparameterization but experience a significant decline in performance. which only tunes continuous prompts with a frozen language model; 3) CMT-PT: It indiscriminately utilizes inter- and intra-sentence context via the concatenation strategy, as depicted in Figure 1 (a). Similar to MT-PT, it is also a tuned LLM via the deep prompt tuning with MLPs reparameterization. Among them, MT-LoRA and MT-PT are context-agnostic systems while CMT-PT is a context-aware system. For a fair comparison, we ensure that all context-aware models built upon LLM, including CMT-PT, MPT, and DeMPT, incorporate identical inter-sentence context. We provide more discussion in utilization of various inter-sentence contexts in Appendix L and M.
Model Setting and Training.
For all encoder-decoder Transformer models, including Transformer (Trans.), MR-Trans and G-Trans101010For G-Trans, we use their official implementation upon Fairseq. Code: https://github.com/baoguangsheng/g-transformer., we implement them upon Fairseq Ott et al. (2019). For MT-LoRA models, we set the rank of trainable matrices as 16 which performs best in our preliminary experiment. For all MT-PT models, CMT-PT models, and our models, we set the prompt length as 64.111111We provide more discussion in Appendix E about the prompt length. For the incorporation of inter-sentence context in CMT-PT and our models, we consider a dynamic , in which the total tokens are no more than 256. In enhanced decoding, we consider the three next word predictions to be equally important by setting both and to 1/3. We provide an analysis of and more training details in Appendix K and B, respectively.
Evaluation.
We use sacreBLEU (accuracy-related metric)121212Signature: nrefs:1|case:mixed|eff:no|tok:13a|
smooth:exp|version:2.3.1 Post (2018), COMET (semantics-related metric) with the wmt22-comet-da model131313https://github.com/Unbabel/COMET Rei et al. (2020), and BlonDe (discourse-related metric) Jiang et al. (2022) as evaluation metrics.141414We provide more discourse-related evaluation in Appendix F.
3.2 Experimental Results
The main experimental results are presented in Tables 1 and 2. Additionally, a comparison of the number of trainable parameters is presented in Table 3 across different tuning methods. When examining llama-2-7b and focusing on context-agnostic models, we find that the Transformer models (Trans.) generally outperform LLMs with LoRA tuning (MT-LoRA) in most translation directions based on BLEU score. However, the MT-LoRA models surpass Trans. in COMET, indicating that translations from LLMs may better align with human preferences. Additionally, the MT-PT models exhibit superior performance compared to the MT-LoRA models across BLEU, COMET, and BlonDe metrics. This improvement could be attributed to the more trainable parameters in the MT-PT models (13.87% vs. 0.12%).
Importantly, by comparing MT-PT and CMT-PT, we observe that CMT-PT which indiscriminately leverages the inter- and intra-sentence context with the concatenation way, even hurts performance for certain translation tasks. For example, the CMT-PT models, despite excelling in discourse-related BlonDe scores (averaging 57.65 vs. 57.17), underperforms in BLEU and COMET compared to the MT-PT models. In contrast, our context-aware MPT and DeMPT models outperform all LLM baselines across all translation tasks in three metrics. For example, our MPT models achieve an average gain of 0.98/0.0103/1.27 in BLEU/COMET/BlonDe compared to the CMT-PT models. Our decoding-enhance strategy further enhances the capacity of LLMs, with DeMPT outperforming MPT with an average gain of 0.65/0.0042/0.88. Compared to G-Trans. (+mBART) or MR-Trans (+mBART), DeMPT also demonstrates either superior or comparable performance across all language pairs.
Finally, we observe a similar performance trend among MT models built upon bloomz-7b1-mt. It also indicates that models built upon llama-2-7b outperform those utilizing bloomz-7b1-mt, suggesting that llama-2-7b serves as a more robust foundation model for translation tasks.
| MT-LoRA | MT-PT/CMT-PT | MPT/DeMPT | |
| Trainable Para. | 0.12% | 13.87% | 3.11% |
4 Discussion
In this section, we use bloomz-7b1-mt as the foundation model to discuss our approach.151515Considering page limitation and the consumption of GPUs resources and training time, we use the ZHEN task as a representative to report the BLEU and BlonDe scores. See Appendix DM for further discussions.
4.1 Effect of Length of Inter-sentence Context
For efficient training, we define the inter-sentence context in Section LABEL:sec:appro as previous sentences with a total tokens not exceeding 256. We are curious about the potential impact of inter-sentence length on the performance of our approach. Consequently, we extend the inter-sentence context length from 256 to 1024 and assess the performance of our approach in the ZHEN task. Figure 4 shows the performance trend of the CMT-PT model and our DeMPT model. As the length of the inter-sentence context increases, both models exhibit a slight enhancement in both BLEU and BlonDe scores. Interestingly, our model with a 256-token inter-sentence context outperforms the CMT-PT model with a 1024-token inter-sentence context in both BLEU and BlonDe scores. This further suggests the effectiveness of our approach in harnessing the capabilities of LLMs for context-aware NMT compared to the concatenation strategy.
4.2 Effect of Multi-phase Strategy
Our multi-phase strategy divides the whole translation into three phases: phase 1 for encoding inter-sentence, phase 2 for encoding intra-sentence, and phase 3 for decoding current sentence. To assess the effect of multi-phase strategy, we compare its performance with two contrasting strategies: merging the first two phases into one (i.e., Merging 1&2) and and merging all phases into a single one (i.e., Merging 1&2&3).161616When merging them all into one, it equals CMT-PT, i.e., the concatenate strategy. Note that in both the two contrasting strategies, we replace the enhanced next word prediction (Eq. 17) in decoding phase with . Table 4 presents the performance of different phrasing strategies. Comparing Merging 1&2 and Merging 1&2&3, it shows that separating the decoding phrase from the encoding marginally improves the performance in both BLEU and BlonDe. Importantly, the comparison of MPT and Merging 1&2 tells that separating the encoding of inter- and intra sentence achieves more gains across all metrics.
| Model | BLEU | COMET | BlonDe |
| MPT | 31.81 | 0.8601 | 50.22 |
| DeMPT | 32.46 | 0.8649 | 50.62 |
[…]
| Merging 1&2 | 31.01 | 0.8503 | 49.91 |
Meanwhile, we conjecture that another benefit of multi-phasing strategy lies in the robustness to the noise contained in document-level context. To test the conjecture, we replace the original inter-sentence context with a random inter-sentence context, meaning we randomly select some sentences from other documents to serve as the inter-sentence context. As shown in Table 5, the performance of both the Merging 1&2&3 and DeMPT models consistently deteriorates when exposed to random context (w/ rnd. CTX). However, the decline is more pronounced for Merging 1&2&3 than for DeMPT (-2.19/0.0102/1.60 vs -0.90/0.0068/0.91). This suggests that DeMPT, owing to its multi-phase strategy, exhibits more robustness in utilizing inter-sentence context in contrast to Merging 1&2&3.
| Model | BLEU | COMET | BlonDe |
| Merging 1&2&3 | 30.82 | 0.8504 | 49.61 |
| w/ rnd. CTX | 28.63 | 0.8402 | 48.01 |
[…]
| w/ rnd. CTX | 31.56 | 0.8581 | 49.71 |
4.3 Human Evaluation
We use the Direct Assessment (DA) method Graham et al. (2017) to manually assess the quality of translations generated by DeMPT and CMT-PT. In this assessment, human evaluators compare the meaning of the MT output with a human-produced reference translation, working within the same language. Specifically, we randomly select 5 documents with a total of 200 groups of sentences from the ZHEN test set. To avoid potential bias in evaluation, we recruit 6 professional translators and ensure each translation from DeMPT or CMT-PT is scored twice by two translators. Table 6 shows the DA scores for CMT-PT and DeMPT. Our DeMPT outperforms CMT-PT by 7.14 DA score, providing strong evidence for the effectiveness of our approach. Further details and results regarding the DA can be found in Appendix H.
| Model | Score_1 | Score_2 | Average |
| CMT-PT | 79.00 | 80.17 | 79.59 |
| DeMPT | 86.17 (+7.17) | 87.30 (+7.13) | 86.73 (+7.14) |
5 Related Work
Due to limited space, we omit the discussion on conventional context-aware MT, focusing instead on LLM-based context-aware MT and prompt tuning for LLMs. Besides, considering our DeMPT’s inspiration from MSP Tan et al. (2022), we offer further discussion on their differences.
LLM-based Context-aware Machine Translation.
While traditional context-aware neural machine translation (NMT) has seen considerable progress in recent years Jean et al. (2017); Wang et al. (2017); Voita et al. (2018); Maruf et al. (2019); Kang et al. (2020); Bao et al. (2021); Sun et al. (2022); Bao et al. (2023), the effective integration of large language models (LLMs) to model inter-sentence context and enhance context-aware translation remains an area of limited exploration. Existing studies mainly focus on the assessment of LLMs’ ability in discourse modeling. For example, Wang et al. (2023) approach context-aware NMT as a task involving long sequence generation, employing a concatenation strategy, and conduct comprehensive evaluations of LLMs such as ChatGPT and GPT-4. Their focus includes the impact of context-aware prompts, comparisons with translation models, and an in-depth analysis of discourse modeling ability. Similarly, Karpinska and Iyyer (2023) engage professional translators to evaluate LLMs’ capacity in context-aware NMT. In contrast, Wu et al. (2024) compare the effectiveness of various parameter-efficient fine-tuning methods on moderately-sized LLMs for context-aware NMT. Besides, Wu and Hu (2023) explore the prompt engineering with GPT language models specifically for document-level (context-aware) MT while Li et al. (2024) experiment with combining sentence-level and document-level translation instructions of varying lengths to fine-tune LLMs. Differently, koneru_etal_2024_contextual propose a post-editing approach to enhance LLMs’ capacity in utilization of inter-sentence context in document-level MT.
Prompt Tuning for Large Language Model.
Liu et al. (2021) and Li and Liang (2021) propose to make LLMs adapt to various tasks by adding trainable prompts (also called continuous prompts) to the original input sequences. In this paradigm, only the continuous prompts are updated during training. Liu et al. (2022) further introduces deep prompt tuning, extending the idea by inserting trainable prompts into all layers of LLMs, rather than just the embedding layer. While these approaches provide a general framework, we focus on enhancing LLM performance specifically for inter-sentence context modeling in context-aware NMT.
Discussion with MSP.
Tan et al. (2022) propose a multi-phase tuning approach (MSP) to enhance the sentence-level translation performance of a multilingual GPT. Our DeMPT mainly differs from MSP in the following aspects: 1) DeMPT adopts a phase-aware prompt to enable distinctive modeling for different inputs, namely inter-sentence contexts, intra-sentence contexts, and the target sentence, a feature not present in MSP; 2) DeMPT incorporates a decoding-enhanced strategy to further improve the effectiveness of utilizing different context information, a capability not available in MSP; 3) DeMPT is designed to alleviate discourse problems in context-aware LLM-based machine translation tasks, rather than addressing sentence-level machine translation tasks as in the case of MSP; 4) DeMPT is designed to adapt LLMs rather than smaller pre-trained model used in MSP.
6 Conclusion
In this paper, we have examined the hypothesis that it is crucial to differentially model and leverage inter-sentence context and intra-sentence context when adapting LLMs to context-aware NMT. This stems from our observation that intra-sentence context exhibits a stronger correlation with the target sentence compared to inter-sentence context, owing to its richer parallel semantic information. To this end, we have proposed a novel decoding-enhanced multi-phase prompt tuning (DeMPT) approach to make LLMs aware of the differences between inter- and intra-sentence contexts, and further improve LLMs’ capacity in discourse modeling. We have evaluated our approach using two foundation models and present experimental results across five translation directions. Experimental results and discussions have demonstrated a significant enhancement in the performance of LLMs in context-aware NMT, manifesting as improved translation accuracy and a reduction in discourse-related issues.
Limitations
Owing to resource limitations, our work is restricted to moderate-scale LLMs, specifically those with 7 billion parameters, and a confined window size of inter-sentence context. It is imperative to acknowledge that the results of our research may differ when employing larger models and extended window sizes for inter-sentence contexts. Considering that English text forms the main body of the training data for LLMs, this paper only focuses on the English-centric translation tasks. The results of non-English-centric translation tasks may vary. We acknowledge these limitations and consider them as avenues for future exploration. Besides, following the finding of Bao et al. (2021), we focus solely on the source-side inter-sentence context in this work. We will explore more about the integration of target-side inter-sentence context in the future.
Appendix A Datasets
Splitting, Preprocessing and Statistics of Datasets.
For all translation tasks, we randomly select 80% document pairs from the corpus as the training set. Both the test set and validation set include 150 document pairs each, randomly sampled from the remaining 20% of document pairs in the corpus. Regarding sentence preprocessing across all datasets for LLM-based models, we segment the sentences with the tokenizer from the respective foundation model. No additional preprocessing steps are performed. For encoder-decoder Transformer models, we segment the source and target sentences into sub-words by a BPE model with 30K merged operations sennrich_etal_acl_2016. We provide the detailed statistic in Table 7. Datasets are downloaded from https://data.statmt.org/news-commentary/v18.
| Dataset | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | |
| Training | 8,622 | 342,495 | 7,915 | 310,489 | 8,417 | 333,201 | 9,677 | 378,281 | 7,255 | 272,100 |
| Validation | 150 | 6,061 | 150 | 5,890 | 150 | 5,866 | 150 | 5,782 | 150 | 5,691 |
| Test | 150 | 5,747 | 150 | 5,795 | 150 | 5,967 | 150 | 5,819 | 150 | 5,619 |
Appendix B Training Details
For all encoder-decoder Transformer NMT models, we use the transformer-base setting as in Vaswani et al. (2017), where the learning rate is set to with an inverse-square schedule and warmup steps of , and use Adam optimizer with and . For the other special training settings in G-Trans and MR-Trans, we keep consistent with that provided in their paper. All Transformer NMT models are trained on 4 NVIDIA V100 32GB GPUs with a batch size of 4096. For the models with prompt tuning in Section LABEL:sec:exp, including MT-PT, CMT-PT, MPT and DeMPT models, the length of the trainable prompt is set as 64. During both training and inference, the model generates only the current target sentence, operating in a many-to-one translation mode. For all fine-tuning models in this paper, we set the training epoch to 4, and the warm-up rate to 0.1. We use the log learning rate decay strategy with a maximum learning rate of 5e-5. We collate a mini-batch by counting the total tokens inside the batch and set the batch size as 4096. All fine-tuning models are trained on 4 NVIDIA A800 GPUs with Deespeed Zero 2 offload setting Rajbhandari et al. (2020).171717https://github.com/microsoft/DeepSpeed
Appendix C Traditional Context-aware Models
In this paper, we implement G-Transformer (G-Trans) Bao et al. (2021) and Multi-Resolution Transformer (MR-Trans) Sun et al. (2022) as representatives of traditional context-aware models for comparison. For ease of understanding, we provide a brief introduction to these two models in this section.
G-Transformer.
The transformer model equips a group attention on the lower encoder/decoder layer and a combined attention on the top encoder/decoder layer. For a sentence being translated with its inter-sentence context, the group attention helps maintain locality bias by focusing on intra-sentence context. Meanwhile, the combined attention effectively integrates boundary information, enhancing the translation process with inter-sentence context.
Multi-Resolution Transformer.
The Transformer model does not include any additional modules specifically for modeling inter-sentence context. Instead, it only uses a mixed training set that comprises both sentence-level and document-level instances with varying numbers of sentences. Training on this mixed set allows the Transformer model to handle both sentence-level and document-level translation tasks. In this paper, we implement its Document-to-Sentence variant, which uses all preceding contexts as the source and the current sentence as the target.
Appendix D Comparison of Inference Speed
| Model | Speed | BLEU |
| MT-PT | 0.75 sec/sent. | 30.99 |
| CMT-PT | 0.77 sec/sent. | 30.82 |
| MPT | 0.78 sec/sent. | 31.81 |
| DeMPT | 0.79 sec/sent. | 32.46 |
Table 8 compares the inference speed of different models on ZHEN translation task. Our MPT and DeMPT models, dividing the context-aware NMT process into three separate phases, demonstrates comparable inference speed to the single-phase MT-PT and CMT-PT models, with only a marginal drop of 0.02 seconds per sentence in decoding. This illustrates the efficiency of our approach without introducing significant computational overhead.
Appendix E Effect of Prompt Length
As our approach is implemented based on deep prompt tuning, next we compare the impact of the trainable prompt length for MT-PT, CMT-PT, and our DeMPT. Figure 7 shows the performance curves when increasing the prompt length from 32 to 128. We observe that increased prompt length tends to enhance performance for both BLEU and BlonDe, yet the gains exhibit diminishing returns. This finding is consistent with that in Li and Liang (2021); Lester et al. (2021); Tan et al. (2022). We also observe that DeMPT with a prompt length of 64 outperforms both MT-PT and CMT-PT with a prompt length of 128 on both metrics, suggesting the superiority of our approach over the concatenation strategy in enhancing LLMs’ capacity for context-aware NMT.
Appendix F Performance on Contrastive Test Set
We evaluate the models’ ability to resolve discourse inconsistencies using the contrastive test set proposed by Voita et al. (2019a), which focuses on four discourse phenomena such as deixis, lexicon consistency (lex.c), ellipsis inflection (ell.infl), and verb phrase ellipsis (ell.VP) in EnglishRussian translation. Within the test set, each instance comprises a positive translation and several negative ones that vary by only one specific word. The purpose of the contrastive test set is to assess whether a model is more inclined to generate a correct translation as opposed to incorrect variations. Table 9 lists the accuracy of translation prediction on the contrastive test set for MT-PT, CMT-PT and DeMPT. Compared to the context-agnostic MT-PT model, both context-aware CMT-PT and DeMPT models show substantial improvements across the four discourse phenomena. Additionally, DeMPT demonstrates the best performance, surpassing CMT-PT by an average accuracy margin of 3.8.
| Model | deixis | lex.c | ell.infl | ell.VP | Avg. |
| MT-PT | 50.0 | 45.7 | 53.0 | 28.6 | 44.3 |
| CMT-PT | 80.2 | 46.1 | 74.3 | 75.3 | 68.9 |
| DeMPT | 80.1 | 55.7 | 75.9 | 79.3 | 72.7 |
Appendix G Effect of Various Contexts for Decoding-enhanced Strategy
| Model | BLEU | COMET | BlonDe |
| MT-PT | 30.99 | 0.8520 | 49.48 |
| CMT-PT | 30.82 | 0.8504 | 49.61 |
[…]
| w/o | 32.33 | 0.8629 | 50.29 |
| w/o | 32.11 | 0.8641 | 50.51 |
We conduct an ablation study on the ZH-EN translation direction using the bloomz-7b-mt model as the foundation model to clarify the effect of the three probabilities in Equation 17, i.e., the effect of various contexts for the heuristic decoding-enhanced strategy. From the Table 10, we observe that removing , i.e., w/o , leads to a significant degradation in the discourse-related metric, namely the BlonDe. This is because the integration enhances the utilization of the inter-sentence context during the decoding phase. We are additionally, removing results in the most substantial degeneration in BLEU metric. This observation demonstrates that our heuristic decoding-enhanced strategy can distinctively improve the utilization of various contexts during the decoding phase.
Appendix H Details of Human Evaluation
Criterion and Recruitment.
Given a source sentence, its translation from MT (i.e., CMT-PT and our DeMPT), and its human-produced reference translation, the evaluators are asked to give a score ranging from 0 to 100. Figure 5 presents the detailed criterion of scoring. We recruit evaluators from professional translators with at least five years of experience in translation.
Statistics of Translation Errors.
We manually count the number of bad cases from our DeMPT model. The bad cases fall into two categories: (1) the DA score is 60 or lower; (2) the DA score is lower than that of the translation from CMT-PT. The main types of the bad cases are Mistranslation (Mis.), Unnoticed Omission (UO), Inappropriate Expression (IE), and Grammatical Error (GE). We present detailed statistics in Table 11. The statistics indicate the bad cases mainly come from Mistranslation and Unnoticed Omission. Meanwhile, our DeMPT model outperforms the CMT-PT model in 86.5% DA cases.
Case Study.
We present a case in Figure 6 to illustrate how our DeMPT model outperforms the CMT-PT model. In this case, we compare the translations of two consecutive sentences from our model and the CMT-PT model. First, we notice that the CMT-PT model translates the source word 美国 in the two sentences into US and America, respectively. However, our model consistently translates them into US. Second, our model uses for its part, a phase with more coherent preference, as the translation of 同时 , instead of At the same time adopted in the translation from the CMT-PT model. Both of them demonstrate the superiority of our proposed approach in discourse modeling.
| Group | Type of Bad Case | ||||
| Mis. | UO | IE | GE | Total (Perc.) | |
| 1 | 6 | 3 | 1 | 2 | 12 (6.0%) |
| 2 | 9 | 7 | 6 | 5 | 27 (13.5%) |
Appendix I Effect of Dataset Scales
| Model | BLEU | COMET | BlonDe |
| CMT-PT | 30.82 | 0.8504 | 49.61 |
| + 200K | 31.21 | 0.8521 | 49.88 |
| + 400K | 31.73 | 0.8555 | 50.11 |
| + 700K | 31.89 | 0.8559 | 50.23 |
[…]
| + 200K | 32.77 | 0.8663 | 50.99 |
| + 400K | 33.56 | 0.8701 | 51.47 |
| + 700K | 33.91 | 0.8721 | 51.97 |
We conduct an experiment to analyze the impact of training dataset scales on the concatenating strategy (CMT-PT) and the multi-phased, decoding-enhanced strategy (DeMPT). To do this, we expand the ZHEN training set with additional document-level data from the LDC.181818The training data set consists of LDC2002T01, LDC2004T07, LDC2005T06, LDC2005T10, LDC2009T02, LDC2009T15, and LDC2010T03. Specifically, we selected 200K,400K and 700K sentence pairs with their inter-sentence context from the LDC and combined them with the existing ZHEN training set to train the CMT-PT and DeMPT models.
Table 12 lists the performances of CMT-PT and DeMPT when extending scales of the training set into 500K (300K +200K), 700K (300k + 400K) and 1M (300K + 700K). We observe increasing the scale of the training set consistently boosts the performance of DeMPT and CMT-PT. However, our DeMPT significantly outperforms CMT-PT across all three metrics.
Appendix J Effect of Transfer Layer and Type Embedding
As in Eq. 22 within Section 2.3, we introduce two sublayers: a non-linear transfer sublayer and a type embedding sublayer for the trainable prompt in each phase. This design enhances the awareness of LLMs regarding the distinctions in inputs across the three tuning phases, allowing them to adapt to specific roles at each phase. We investigate the effect of these two sublayers.
As shown in Table 15, our observations reveal that the transfer sublayer holds greater importance than the type embedding sublayer. Removing either the non-linear transfer sublayer (w/o Transfer.) or the type embedding sublayer (w/o Embed.) results in a performance drop of 0.84/0.0048/0.39 or 0.45/0.0036/0.007 in BLEU/COMET/BlonDe metrics.
| Model | BLEU | COMET | BlonDe |
| MT-PT | 30.99 | 0.8520 | 49.48 |
| CMT-PT | 30.82 | 0.8504 | 49.61 |
[…]
| w/o Transfer. | 31.62 | 0.8601 | 50.23 |
| w/o Embed. | 32.01 | 0.8613 | 50.55 |
| w/o CTX. | 31.98 | 0.8593 | 49.89 |
Appendix K Effect of Hyperparameter
| Model (DeMPT) | BLEU | COMET | BlonDe |
| =1/3, =1/3, ZHEN | 32.46 | 0.8649 | 50.62 |
| =1/4, =1/3, ZHEN | 32.51 | 0.8653 | 50.31 |
[…]
| =1/4, =1/3, FREN | 41.82 | 0.8785 | 57.92 |
Due to the limited computational resources, we do not perform extensive experiments to find the optimal combination of and for different translation tasks, simply setting them to be equal. For example, verifying each combination of and requires 10 experiments (5 2 for the number of translation directions and foundation models). Therefore, we carry out targeted experiments using a combination of and on ZHEN and FREN only here.
The results are reported in Table 14. We use a smaller value for here and observe that the BlonDe scores are more sensitive to changes compared to BLEU and COMET. For example, a smaller results in -0.31 and -0.38 for ZHEN and FREN, respectively. This sensitivity may be reasonable because is used for adjusting the utilization of inter-sentence context.
Appendix L Effect of Inter-sentence Context
We implement the context-agnostic (sentence-level) DeMPT system to analyze the effect of the inter-sentence context and differences with MSP. More specifically, we replace the input of LLMs in the inter-sentence context encoding phase with the intra-sentence context. In other words, we encode the intra-sentence context twice to keep the multi-phase tuning strategy in DeMPT while making the inter-sentence context unavailable.
As shown in the last row of Table 15 (i.e., w/o CTX), we find that the inter-sentence context is crucial for the alleviation of discourse-related issues. The BlonDe score drops by 0.73 when the inter-sentence context is unavailable. Meanwhile, our DeMPT also significantly improves the performance of LLMs in context-agnostic MT, e.g., + 0.99 BLEU score and + 0.0073 COMET score compared to the MT-PT model.
Appendix M Effect of Target-side Inter-sentence Context
| Model | d-BLEU | d-COMET | d-BlonDe |
| MT-PT (m2o) | 34.19 | 0.8216 | 49.48 |
| CMT-PT (m2o) | 34.06 | 0.8211 | 54.68 |
| DeMPT (m2o) | 35.76 | 0.8316 | 55.97 |
[…]
To enable a fair comparison, we incorporate only the source-side inter-sentence context for the model with the concatenating strategy, i.e., the CMT-PT model in the many-to-one (m2o) translation mode, as shown in Tables 1 and 2. To further investigate the effect of target-side inter-sentence context for the concatenating strategy, we compare the CMT-PT model in the many-to-many (m2m) translation mode to the models in the many-to-one translation mode, for the ZHEN translation task when using the bloomz-7b1-mt as the foundation model.
Different from the results in Tables 1 and 2, we report the document-level BLEU, BlonDe, and COMET scores for all models here due to the unavailability of sentence-level alignment for many-to-many model. From the experimental results, we observe that the CMP-PT (m2m) model outperforms the CMP-PT (m2o) model (mostly significant in terms of the d-BlonDe metric), which demonstrates the effectiveness of the target context in addressing discourse issues. However, the CMP-PT (m2m) model still underperforms the DeMPT model across three metrics.